OpenStudy (anonymous):

solve for \[\left( \sqrt{5x}-3\right)^{2}\]

OpenStudy (anonymous):

You can't solve that unless it equals something

OpenStudy (saifoo.khan):

You can

OpenStudy (anonymous):

I guess you can expand it and then try solving for x

OpenStudy (saifoo.khan):

\[(\sqrt{5x}-3)(\sqrt{5x}-3)\]Solve.

OpenStudy (anonymous):

well i think you mean to make it a polynomial in which case it would be \[(\sqrt5x-3)(\sqrt5x-3)\]\[5x-3\sqrt5-3\sqrt5+9\]\[5x -6\sqrt5 +9\]

OpenStudy (anonymous):

^ yep, that's where I was going too

OpenStudy (anonymous):

but there is nothing to solve for so you can't solve for x

OpenStudy (saifoo.khan):

Agreeeeeee.

OpenStudy (anonymous):

\[9-6 \sqrt{5} \sqrt{x}+5 x\]

OpenStudy (anonymous):

if it was 5x-6(sqrt5) +9 = 0 you could

OpenStudy (anonymous):

\[-6 \sqrt{5} \sqrt{x}+5 x=-9\] x=5/9

OpenStudy (anonymous):

but that wasn't part of the problem, imran. you can't do that.

OpenStudy (anonymous):

its a problem not an equation. there is no other side that you can subtract the 9 to.